Mandy's Candies is most famous for their homemade fudge and homemade toffee. They make an average of 154 pounds of fudge and toffee combined each month. They sell a pound of fudge for $9.64 and a pound of toffee for $6.44. Mandy's averages $1,231.76 each month in fudge and toffee sales. On average, how many pounds each of fudge and toffee does Mandy's Candies sell in one month?
Found 2 solutions by mananth, mbmodak:
Answer by mananth(16946)

x= number of fudges
y= number of toffees
Total sold
1.00 x + 1.00 y = 154.00 .............1
Total value
9.64 x + 6.44 y = 1231.76 .............2
Eliminate y
multiply (1)by -6.44
Multiply (2) by 1.00
-6.44 x -6.44 y = -991.76
9.64 x + 6.44 y = 1231.76
Add the two equations
3.20 x = 240.00
/ 3.20
x = 75.00
plug value of x in (1)
1.00 x + 1.00 y = 154.00
75.00 + y = 154.00
y = 154.00 -75.00
y = 79.00
y = 79.00
x= 75.00 number of fudges
y= 79.00 number of toffees
m.ananth@hotmail.ca

Answer by mbmodak(32)

Let qty of fudge be x pounds & qty of toffee be y pounds per month

They make an average of 154 pounds of fudge and toffee combined each month

x+y=154 (1)

They sell a pound of fudge for $9.64 and a pound of toffee for $6.44. Mandy's averages $1,231.76 each month in fudge and toffee sales. On average,

So cost of monthly fudge&toffee is 9.64*x+6.44*y=1231.76

9.64x+6.44y=1231.76 ;(2)

multiply eq (1) by 9.64 & then subtract 2 from 1

9.64x+9.64y=154*9.64
9.64x+9.64y=1484.56 (1')
9.64x+6.44y=1231.76 ;(2)
-----------------------
(9.64-6.44)y=1484.64-1231.76

3.2y=252.8

y=252.8/3.2=79 pounda

In order to solve for x, take the value for y and substitute it back into either one of the original equations.

x+79=154

x=154-79=75 pounds So they make 75 pounds of fudge & 79 pounds of toffee per month
so monthly cost is
9.64*75+6.44*79=1231.76